OpenStudy (anonymous):

how do I evaluate the limit of ( 1 - 1/x)e^1/x as x approaches 0-? I have seen someone taking s = 1/x so then it becomes ( 1-s)e^s but he said that x approaches - infinity. Can anyone explain?

OpenStudy (anonymous):

Means what it is... Given 1/x as x goes to 0, 1/x gets bigger and bigger, means it goes to infinity So when x goes to 0, 1/x, or s, goes to infinity :)

OpenStudy (anonymous):

oh wow thanks for that :)

OpenStudy (anonymous):

No sweat :)

OpenStudy (anonymous):

It's a substitution. \[x = \frac{1}{s}\iff s=\frac{1}{x}, \text{ so as $x\to0^+$, $s\to\infty$}\]